# Transforming Indoor Localization: Advanced Transformer Architecture for NLOS Dominated Wireless Environments with Distributed Sensors

### v1

@@ IV-A Pre-processing for RF Signals from Distributed Sensors
To address this dynamic range issue, we apply the power compression algorithm [13] on the PDP before feeding the signals to the DL model. For a length-$N_{\text{ts}}$ PDP $p[d]$, the power-compressed signal is given by:
@@
where $\|\{p[d]\}\|=\sum_{d=0}^{N_{\text{ts}}-1}p[d]$ is the total received power at a sensor over all its antenna ports, $r$ is the amplitude compression ratio parameter and $S$ is the target scale parameter after compression. For our indoor factory scenario in Fig. 1, we use $r=5$ and $S=10$ to compress the received signal powers into a narrower range between 00 and $10$ dB illustrated in Fig. 2. Furthermore, we found better performance is obtained when using square root PDP, $\sqrt{\check{p}[d]}$, instead of straight PDP.
@@ IV-B Conventional Tokenization Approaches
Patch-based Tokenization (PBT): In the wireless positioning literature, Transformer inputs are often represented in image format, with some approaches treating the CSI matrix as an image [17], and others converting RSSI and AoA data into images [15], [19]. For the first tokenization approach, similar to methods used in wireless literature, the PDP $\Psi$ which is a matrix of dimensions $N_{\rm S}\times N_{\textrm{ts}}$, is treated as a 2D image. This image is partitioned into fixed-size patches (referred to as tokens) with dimensions $W_{\rm h}\times W_{\rm w}$, where $W_{\rm h}$ and $W_{\rm w}$ denote the height and width of each token, respectively. Consequently, the total number of tokens generated from the PDP is given by $N_{\rm tk}=\frac{N_{\rm S}}{W_{\rm h}}\cdot\frac{N_{\text{ts}}}{W_{\rm w}}$. Each token is then reshaped into a flattened vector, resulting in a spatial size (i.e., number of sample per token) of $N_{\rm st}=W_{\rm h}\cdot W_{\rm w}$, this tokenization technique is referred to as Patch-based Tokenization (PBT). However, transforming inputs into image format and applying patching techniques fuses information from distinct physical measurements without considering the physical meaning of samples, which may hinder the Transformer from learning meaningful attention maps without effectively capturing the nuances of wireless communication environments. The PBT framework has not been previously utilized in conjunction with PDP in the literature. To the best of our knowledge, this paper is the first to explore its application with PDP.
@@ IV-C Proposed Tokenization Approaches
(a) Vanilla Transformer model.
(b) Lightweight Swish-Gated Transformer model (L-SwiGLU ViT).
Figure 4: The architecture of the vanilla Transformer model (Fig. 4(a)), while the proposed lightweight Swish-Gated Transformer (Fig. 4(b)), with the modified parts highlighted in red.
@@ V Designing Transformer Model for Distributed Sensor Networks
With the recent success of encoder-only Transformer architectures like ViT in computer vision, there has been growing interest in adapting them for indoor positioning. However, directly applying ViT to this domain results in suboptimal performance and significantly increases model complexity. In this paper, we propose leveraging an encoder-only Transformer model for device localization in cluttered indoor environments characterized by highly NLOS conditions. Specifically, the Transformer is designed to learn the complex mapping $\mathsf{f}$ between the device’s position $\hat{\bm{y}}=[x_{\rm d},y_{\rm d}]$ and the PDP $\mathbf{\Psi}$ received from all sensor nodes $N_{\rm S}$, as expressed by:
$\hat{\bm{y}}=\mathsf{f}(\bm{P},\mathbf{\Psi};\Theta)~{},$ (6)
where $\Theta$ encapsulates the model parameters, and $\bm{P}$ is the sensor position vector as defined in Section III. The PDP, obtained from channel measurements at the sensors, encodes the multipath propagation characteristics by capturing the power and delay of reflected and scattered signals. This information inherently represents the spatial geometry and structure of the environment, enabling the Transformer to accurately learn the relationship between the PDP and the device’s position $\hat{\bm{y}}$.
In this section, we describe the architecture of the vanilla Transformer model (i.e., ViT) and the proposed lightweight Swish-Gated Linear Unit-based Transformer model, referred to as L-SwiGLU Transformer. The complete architecture of vanilla Transformer and L-SwiGLU Transformer model can be seen in the Fig. 4.
V-A Vanilla Transformer Architecture
Fig. 4(a) illustrates the architecture of the vanilla Transformer model. The localization process using a Transformer can be categorized into $3$ main steps: $1$) Input Embedding Preparation $2$) Encoder Mechanism, and $3$) Position Estimation.
@@ V-A2 Encoder Mechanism
where $d_{k}=D_{\text{emb}}$ serves as the normalization term. The resulting correlation matrix $\mathbf{C}$ quantifies the similarity between tokens, with each element representing the degree of correlation between corresponding token pairs. Then, the softmax is applied along the rows to turn similarity values into probabilities.
@@
where $j\in\{1,2,\cdots,N_{L}\}$. This operation assigns a weighted sum of the values ($\mathbf{V}$) to each embedding, where the weights are determined by the similarity between the query ($\mathbf{Q}$) and key ($\mathbf{K}$) vectors. The similarity calculation is essential, as it enables the model to learn relationships between tokens by distributing attention based on relevance, rather than solely relying on exact matches between $\mathbf{Q}$ and $\mathbf{K}$.
In [7], the authors hypothesized that instead of projecting the embeddings to generate $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$ one time, it is more beneficial to project it $N_{\rm h}$ times using $N_{\rm h}$ different learnable weight matrices $\mathbf{W}^{i}_{q}$, $\mathbf{W}^{i}_{k}$, and $\mathbf{W}^{i}_{v}$, $i\in(1,\cdots,N_{\rm h})$, where $N_{\rm h}$ denotes the total number of heads. In this way, the model can learn $N_{\rm h}$ different representations from the input embeddings. To maintain the computational complexity of the multi-head attention, block similar to that of the single-head block, one can choose $d_{q}^{i}=d_{k}/N_{\rm h}$, with the same selection for $d_{k}^{i}$ and $d_{v}^{i}$. The overall complexity of $\mathsf{MHA}$ attention layer is $O(\tilde{N}_{\rm tk}^{2}D_{\text{emb}})$.
@@
where $\sigma$ is the activation function and $\mathbf{W}^{1}_{j},\mathbf{W}^{2}_{j},b^{1}_{j},$ and $b^{2}_{j}$ are the weight and biases of $\mathsf{MLP}$ layers in $j^{\rm th}$ Transformer block.
@@ V-B Proposed L-SwiGLU ViT
This subsection provides a comprehensive step-by-step explanation of the modifications made to the Transformer components. The Fig. 4(b) depicts the architecture of the proposed L-SwiGLU Transformer model, highlighting the introduced enhancements. The modifications made in the proposed model are categorized into two parts: $1$) Modifications Within the Transformer Encoder Block, and $2$) Modifications Outside the Transformer Encoder Block. These modifications are detailed in the following subsections.
@@ V-B1 Modifications Within the Transformer Encoder Block
To enhance the stability, robustness, and overall performance of the L-SwiGLU model, the standard $\mathsf{LN}$ used in the vanilla Transformer model is replaced with a Root Mean Square (RMS) normalization layer. LayerNorm performs both re-centering and re-scaling, making it robust to shift noise in inputs and weights. However, in line with the hypothesis proposed in [22]—which suggests that the success of $\mathsf{LN}$ is attributed primarily to its re-scaling invariance rather than re-centering invariance—we substitute $\mathsf{LN}$ with $\mathsf{RMSNorm}$. $\mathsf{RMSNorm}$ exclusively preserves re-scaling invariance while omitting re-centering, aligning with this hypothesis to further optimize the model’s performance. The $\mathsf{RMSNorm}$ operation for the input $\tilde{\bm{Z}}_{\rm j-1}$ is defined as:
@@
In the proposed L-SwiGLU model, the standard $\mathsf{MLP}$ is replaced with a Swish-Gated Linear Unit (SwiGLU)-based $\mathsf{MLP}$ to enhance performance [23]. Recognizing that not all information within each token contributes equally to predicting the location, we incorporated the Gated Linear Unit (GLU) within the $\mathsf{MLP}$ block to enable the model to focus on the most relevant features. Acting as an effective denoiser, the GLU is particularly beneficial in wireless communication scenarios where filtering out irrelevant data is critical. The GLU achieves this by performing a component-wise product of two linear projections of size $h^{\text{SwiGLU}}_{\text{dim}}$, selectively passing meaningful information.
@@
where the output of the $\mathsf{MHA}$ block, $\tilde{\bm{Z}}_{j}^{\mathsf{MHA}}$, is first normalized using $\mathsf{RMSNorm}$ and then processed by the $\mathsf{SwiGLU}$ block (16). The resulting transformed representation is combined with the residual $\mathsf{MHA}$ output $\tilde{\bm{Z}}_{j}^{\mathsf{MHA}}$ (11), thereby enhancing representational capacity while preserving gradient stability.
@@ V-B2 Modifications Outside the Transformer Encoder Block
In the vanilla ViT model, a class token $\mathbf{X}_{\text{cls}}$ is appended to the input sequence to capture the global representation, which is subsequently used for prediction. However, the inclusion of the class token introduces an additional token to those generated during tokenization from PDP. Since the $\mathsf{MHA}$ layer exhibits quadratic complexity with respect to the number of tokens ($\tilde{N}_{\rm tk}$), this addition significantly increases the computational complexity. To address this issue, the class token mechanism is replaced with a Global Average Pooling (GAP) layer. The GAP layer, placed after the $N_{\rm L}$ Transformer encoder, replaces the class token mechanism by aggregating information from all tokens through averaging across the token dimension. This approach enables the model to utilize the collective representation for prediction while significantly reducing computational overhead. In this case, the total number of tokens in the embedding $\mathbf{X}_{\text{emb}}=[\mathbf{X}_{\text{inp}}]\in\mathbb{R}^{\tilde{N}_{\rm tk}\times D_{\text{emb}}}$ becomes $\tilde{N}_{\rm tk}=N_{\rm tk}$. The RMS normalization is applied on top of the aggregated information from the GAP layer, ensuring a well-normalized representation. This normalized output is then passed to the $\mathsf{IH}$ for the final prediction:
@@
Furthermore, due to the proposed tokenization approach described in Section V, positional embeddings are no longer required. In this approach, the entire sequence of sensor data is treated as a single token, and since the sensor positions are fixed, the positional order is inherently stored within the token itself. As a result, the positional embedding layer used in the vanilla Transformer can be omitted, simplifying the model architecture and reducing computational overhead:
@@
An additional advantage of eliminating positional embeddings is that it prevents the introduction of bias into token representations, which can disrupt the modeling of multivariate correlations. By removing positional embeddings, the proposed model retains the integrity of token information, enhancing the $\mathsf{MHA}$ block’s capability to accurately capture complex multivariate dependencies.
@@ VI Evaluation and Analysis
For wireless sensor applications, computation and energy resources are typically highly constrained. In this study, to evaluate the performance of the proposed model and tokenization approach, we consider small, medium, and large transformer models, each with strict limits on the number of FLOPs: 4.5M, 16.5M, and 63.5M FLOPs, respectively. These constraints enable a comprehensive analysis of how data requirements and computational demands vary across models of different sizes.
The remaining part of this section is structured as follows. First, we outline the training approach. Next, we provide a comparison of the proposed tokenization methods with the vanilla architecture. Finally, we evaluate the proposed L-SwiGLU architecture, utilizing the SST method, against the vanilla transformer.
@@ VI-A Training Approach
The availability of large datasets has been a major contributor to the success of DL models. However, collecting large quantity of RF data and the corresponding ground truth device locations is very costly and time-consuming. There is hence a strong need for effective data augmentation techniques to alleviate the data collection burden. We adopt the three RF data augmentation techniques introduced in [13] and summarize them briefly in the following.
@@ VI-A2 Random Signal Shifting
Each of the 18 TD PDPs is independently shifted by a random amount along the time axis:
@@
where $\delta$ follows a zero-mean truncated normal distribution with parameter $\sigma_{\text{RSS}}^{2}$ and limits of $(-2\sigma_{\text{RSS}},2\sigma_{\text{RSS}})$. This mimics the effect of network synchronization or transmit/receive timing offsets between the sensors and device, which are common impairments in real-world scenarios. We follow the suggested setting of $\sigma_{\text{RSS}}=25$ ns.
@@ VI-A3 Smoothed Regression Mixup (SRM)
This is an improved mixup approach for regression problems. Two data samples, A and B, with the dimension $18\times 256$ PDPs, $x_{A}$ and $x_{B}$, are selected from a mini-batch to synthesize a new mixup PDP: $x_{C}=\lambda x_{A}+(1-\lambda)x_{B}$, based on the probability given by a Gaussian kernel $\exp\left(-\frac{d(A,B)}{2\sigma_{\text{mix}}^{2}}\right)$ [24]. Here, $\lambda$ is a beta distributed random mixing ratio with parameter (2,2), $d(A,B)$ is a similarity measure between the two data samples, and $\sigma_{\text{mix}}^{2}$ is a hyperparameter. The probability kernel is designed such that two samples more similar to each other are more likely used as mixup inputs, while two rather different samples are rarely paired for mixing. Given the mixup PDP input of $x_{C}$, the loss function for the model output $\hat{y}_{C}$ is computed as the weighted average of the $L_{1}$ losses against the pre-mixing labels:
@@
(a) Small dataset.
(b) Medium dataset.
(c) Large dataset.
Figure 7: Comparison of the CDF of 2D positioning errors for large models on different dataset sizes, evaluated across three tokenization methods using the Vanilla Transformer.
(a) Small dataset.
(b) Medium dataset.
(c) Large dataset.
Figure 8: Comparison of $\Delta_{90}$ positioning errors for varying model and dataset sizes, evaluated across three tokenization methods using the Vanilla Transformer.
@@
where, $\theta^{\text{EMA}}_{t}$ and $\theta^{\text{EMA}}_{t-1}$ represent the EMA at epoch $t$ and $t-1$, respectively, while $\theta_{t}$ denotes the model weight at epoch $t$.

### v2

@@ IV-A Pre-processing for RF Signals from Distributed Sensors
To address this dynamic range issue, we apply the power compression algorithm [khamesi2024] on the PDP before feeding the signals to the DL model. For a length-$N_{\text{ts}}$ PDP $p[d]$, the power-compressed signal is given by:
@@
where $\|\{p[d]\}\|=\sum_{d=0}^{N_{\text{ts}}-1}p[d]$ is the total received power at a sensor over all its antenna ports, $r$ is the amplitude compression ratio parameter and $S$ is the target scale parameter after compression. For our indoor factory scenario in Fig. 1, we use $r=5$ and $S=10$ to compress the received signal powers into a narrower range between 0 and $10$ dB illustrated in Fig. 2. Furthermore, we found better performance is obtained when using square root PDP, $\sqrt{\check{p}[d]}$, instead of straight PDP.
@@ IV-B Conventional Tokenization Approaches
Patch-based Tokenization (PBT): In the wireless positioning literature, Transformer inputs are often represented in image format, with some approaches treating the CSI matrix as an image [prasad2024vision], and others converting RSSI and AoA data into images [zhou2024vtil], [pan2022transformer]. For the first tokenization approach, similar to methods used in wireless literature, the PDP $\Psi$ which is a matrix of dimensions $N_{\rm S}\times N_{\textrm{ts}}$, is treated as a 2D image. This image is partitioned into fixed-size patches (referred to as tokens) with dimensions $W_{\rm h}\times W_{\rm w}$, where $W_{\rm h}$ and $W_{\rm w}$ denote the height and width of each token, respectively. Consequently, the total number of tokens generated from the PDP is given by $N_{\rm tk}=\frac{N_{\rm S}}{W_{\rm h}}\cdot\frac{N_{\text{ts}}}{W_{\rm w}}$. Each token is then reshaped into a flattened vector, resulting in a spatial size (i.e., number of sample per token) of $N_{\rm st}=W_{\rm h}\cdot W_{\rm w}$, this tokenization technique is referred to as Patch-based Tokenization (PBT). However, transforming inputs into image format and applying patching techniques fuses information from distinct physical measurements without considering the physical meaning of samples, which may hinder the Transformer from learning meaningful attention maps without effectively capturing the nuances of wireless communication environments. The PBT framework has not been previously utilized in conjunction with PDP in the literature. To the best of our knowledge, this paper is the first to explore its application with PDP.
@@ IV-C Proposed Tokenization Approaches
((a))
((b))
Figure 4: The architecture of the vanilla-T model, an encoder-only Transformer, is depicted in Fig. 4(a), while the proposed lightweight Swish-Gated Transformer, with the modified parts highlighted in light red, is shown in Fig. 4(b).
@@ V Designing Transformer Model for Distributed Sensor Networks
In this work, we design an encoder-only Transformer architecture tailored for device localization in cluttered indoor environments dominated by severe NLoS conditions. Specifically, the Transformer is designed to learn the complex mapping $f_{\bm{\Theta}}$ between the device’s position $y$ and the PDP $\mathbf{\Psi}$ received from all sensor nodes $N_{\rm S}$. The PDP, obtained from channel measurements at the sensors, encodes the multipath propagation characteristics by capturing the power and delay of reflected and scattered signals. This information inherently represents the spatial geometry and structure of the environment, enabling the Transformer to accurately learn the relationship between the PDP and the device’s position $\hat{\bm{y}}$.
In this section, we describe the architectures of the baseline Vanilla Transformer (Vanilla-T) and the proposed Lightweight Swish-Gated Linear Unit Transformer (L-SwiGLU-T). The complete model structures are illustrated in Fig. 4.
V-A Vanilla Transformer (Vanilla-T) Architecture
Fig. 4(a) illustrates the architecture of the Vanilla-T model. The localization process using a Transformer can be categorized into $3$ main steps: $1$) Input Embedding Preparation $2$) Encoder Mechanism, and $3$) Position Estimation.
@@ V-A2 Encoder Mechanism
where $d_{k}=D_{\text{emb}}$ serves as the normalization term. The $\mathbf{C}$ quantifies the similarity between tokens, with each element representing the degree of correlation between corresponding token pairs. Then, the softmax is applied along the rows to turn similarity values into probabilities.
@@
where $j\in\{1,2,\cdots,N_{L}\}$. This operation assigns a weighted sum of the values ($\mathbf{V}$) to each embedding, where the weights are determined by the similarity between the query ($\mathbf{Q}$) and key ($\mathbf{K}$) vectors. The similarity calculation enables the model to learn relationships between tokens by distributing attention based on relevance, rather than solely relying on exact matches between $\mathbf{Q}$ and $\mathbf{K}$.
In [dosovitskiy2020image], the authors hypothesized that instead of projecting the embeddings to generate $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$ one time, it is more beneficial to project it $N_{\rm h}$ times using $N_{\rm h}$ different learnable weight matrices $\mathbf{W}^{i}_{q}$, $\mathbf{W}^{i}_{k}$, and $\mathbf{W}^{i}_{v}$, $i\in(1,\cdots,N_{\rm h})$, where $N_{\rm h}$ denotes the total number of heads. In this way, the model can learn $N_{\rm h}$ different representations from the input embeddings. To maintain the computational complexity of the multi-head attention, block similar to that of the single-head block, one can choose $d_{q}^{i}=d_{k}/N_{\rm h}$, with the same selection for $d_{k}^{i}$ and $d_{v}^{i}$. The overall complexity of $\mathsf{MHA}$ attention layer is $O(\tilde{N}_{\rm tk}^{2}D_{\text{emb}})$.
@@
where $\sigma$ is activation function and $\mathbf{W}^{1}_{j},\mathbf{W}^{2}_{j},b^{1}_{j},$ and $b^{2}_{j}$ are weight and biases of $\mathsf{MLP}$ layers in $j^{\rm th}$ Transformer block.
@@ V-B Proposed L-SwiGLU ViT
This subsection provides a comprehensive step-by-step explanation of the modifications made to the Transformer components. The Fig. 4(b) depicts the architecture of the proposed L-SwiGLU-T model, highlighting the introduced enhancements. The modifications made in the proposed model are categorized into two parts: $1$) Modifications Within the Transformer Encoder Block, and $2$) Modifications Outside the Transformer Encoder Block. These modifications are detailed in the following subsections.
@@ V-B1 Modifications Within the Transformer Encoder Block
To enhance the stability, robustness, and overall performance of the L-SwiGLU-T model, the standard $\mathsf{LN}$ used in the vanilla-T model is replaced with a Root Mean Square (RMS) normalization layer. LayerNorm performs both re-centering and re-scaling, making it robust to shift noise in inputs and weights. However, in line with the hypothesis proposed in [zhang2019root]—which suggests that the success of $\mathsf{LN}$ is attributed primarily to its re-scaling invariance rather than re-centering invariance—we substitute $\mathsf{LN}$ with $\mathsf{RMSNorm}$. $\mathsf{RMSNorm}$ exclusively preserves re-scaling invariance while omitting re-centering, aligning with this hypothesis to further optimize the model’s performance. The $\mathsf{RMSNorm}$ operation for the input $\tilde{\bm{Z}}_{\rm j-1}$ is defined as:
@@
In the proposed L-SwiGLU-T model, the standard $\mathsf{MLP}$ is replaced with a Swish-Gated Linear Unit (SwiGLU)-based $\mathsf{MLP}$ to enhance performance [shazeer2020glu]. Recognizing that not all information within each token contributes equally to predicting the location, we incorporated the Gated Linear Unit (GLU) within the $\mathsf{MLP}$ block to enable the model to focus on the most relevant features. Acting as an effective denoiser, the GLU is particularly beneficial in wireless communication scenarios where filtering out irrelevant data is critical. The GLU achieves this by performing a component-wise product of two linear projections of size $h^{\text{SwiGLU}}_{\text{dim}}$, selectively passing meaningful information.
@@
where the output of the $\mathsf{MHA}$ block, $\tilde{\bm{Z}}_{j}^{\mathsf{MHA}}$, is first normalized using $\mathsf{RMSNorm}$ and then processed by the $\mathsf{SwiGLU}$ block (16). The transformed representation is combined with the residual $\mathsf{MHA}$ output $\tilde{\bm{Z}}_{j}^{\mathsf{MHA}}$ (11), thereby enhancing representational capacity while preserving gradient stability.
@@ V-B2 Modifications Outside the Transformer Encoder Block
In the vanilla-T model, a class token $\mathbf{X}_{\text{cls}}$ is appended to the input sequence to capture the global representation, which is subsequently used for prediction. However, the inclusion of the class token introduces an additional token to those generated during tokenization from PDP. Since the $\mathsf{MHA}$ layer exhibits quadratic complexity with respect to the number of tokens ($\tilde{N}_{\rm tk}$), this addition significantly increases the computational complexity. To address this issue, the class token mechanism is replaced with a Global Average Pooling (GAP) layer. The GAP layer, placed after the $N_{\rm L}$ Transformer encoder, replaces the class token mechanism by aggregating information from all tokens through averaging across the token dimension. This approach enables the model to utilize the collective representation for prediction while significantly reducing computational overhead. In this case, the total number of tokens in the embedding $\mathbf{X}_{\text{emb}}=[\mathbf{X}_{\text{inp}}]\in\mathbb{R}^{\tilde{N}_{\rm tk}\times D_{\text{emb}}}$ becomes $\tilde{N}_{\rm tk}=N_{\rm tk}$. The RMS normalization is applied on top of the aggregated information from the GAP layer, ensuring a well-normalized representation. This normalized output is then passed to the $\mathsf{IH}$ for the final prediction:
@@
Furthermore, due to the proposed tokenization approach described in Section V, positional embeddings are no longer required. In this approach, the entire sequence of sensor data is treated as a single token, and since the sensor positions are fixed, the positional order is inherently stored within the token itself. As a result, the positional embedding layer used in the vanilla-T can be omitted, simplifying the model architecture and reducing computational overhead:
@@
An additional advantage of eliminating positional embeddings is that it prevents the introduction of bias into token representations, which can disrupt the modeling of multivariate correlations.
The L-SwiGLU-T architecture integrates RMSNorm, SwiGLU feed-forward layers, GAP, and the removal of positional embeddings to form a domain-optimized Transformer variant for PDP-based indoor localization. Each modification is physically motivated: RMSNorm stabilizes wide dynamic-range inputs, SwiGLU selectively filters relevant multipath components, GAP reduces quadratic attention cost, and eliminating positional embeddings avoids introducing bias while simplifying the model structure.
@@ VI Evaluation and Analysis
For wireless sensor applications, computation and energy resources are typically highly constrained. In this study, to evaluate the performance of the proposed model and tokenization approach, we consider small, medium, and large Transformer models, each with strict limits on the number of FLOPs mathcalBmathcal{B}: 4.5M, 16.5M, and 63.5M FLOPs, respectively. These constraints enable a comprehensive analysis of how data requirements and computational demands vary across models of different sizes.
The remaining part of this section is structured as follows. First, we outline the training approach. Next, we provide a comparison of the proposed tokenization methods with the vanilla architecture. Finally, we evaluate the proposed L-SwiGLU-T architecture, utilizing the SST method, against the vanilla-T.
@@ VI-A Training Approach
The availability of large datasets has been a major contributor to the success of DL models. However, collecting large quantity of RF data and the corresponding ground truth device locations is very costly and time-consuming. There is hence a strong need for effective data augmentation techniques to alleviate the data collection burden. We adopt the three RF data augmentation techniques introduced in [khamesi2024] and summarize them briefly in the following.
@@ VI-A2 Random Signal Shifting
Each of the 18 TD PDPs is independently shifted randomly along the time axis.
@@
where, $\overset{\leftrightarrow}{p}[d]$ represents the delay-shifted PDP and $\delta$ follows a zero-mean truncated normal distribution with parameter $\sigma_{\text{RSS}}^{2}$ and limits of $(-2\sigma_{\text{RSS}},2\sigma_{\text{RSS}})$. This mimics the effect of network synchronization or transmit/receive timing offsets between the sensors and device, which are common impairments in real-world scenarios. We follow the suggested setting of $\sigma_{\text{RSS}}=25$ ns.
@@ VI-A3 Smoothed Regression Mixup (SRM)
This is an improved mixup approach for regression problems. Two data samples, A and B, with the dimension $18\times 256$ PDPs, $x_{A}$ and $x_{B}$, are selected from a mini-batch to synthesize a new mixup PDP: $x_{C}=\lambda x_{A}+(1-\lambda)x_{B}$, based on the probability given by a Gaussian kernel $\exp\left(-\frac{d(A,B)}{2\sigma_{\text{mix}}^{2}}\right)$ [yao2022c]. Here, $\lambda$ is a beta distributed random mixing ratio with parameter (2,2), $d(A,B)$ is a similarity measure between the two data samples, and $\sigma_{\text{mix}}^{2}$ is a hyperparameter. The probability kernel is designed such that two samples more similar to each other are more likely used as mixup inputs, while two rather different samples are rarely paired for mixing. Given the mixup PDP input of $x_{C}$, the loss function for the model output $\hat{y}_{C}$ is computed as the weighted average of the $L_{1}$ losses against the pre-mixing labels:
@@
((a))
((b))
((c))
Figure 7: Comparison of the CDF of 2D positioning errors for large models on different dataset sizes, evaluated across three tokenization methods using the Vanilla-T.
((a))
((b))
((c))
Figure 8: Comparison of $\Delta_{90}$ positioning errors for varying model and dataset sizes, evaluated across three tokenization methods using the Vanilla-T.
@@
where, $\theta^{\text{EMA}}_{t}$ and $\theta^{\text{EMA}}_{t-1}$ represent the EMA at epoch $t$ and t1t-1, while $\theta_{t}$ denotes the model weight at epoch $t$.